OpenStudy (anonymous):

Help me! Medal will be given! What are the solution(s) to the following system of equations? y = x2 y = –x Anwser: (–1, 1) and (1, 1) (0, 0) (0, 0) and (–1, 0) (0, 0) and (–1, 1)

OpenStudy (anonymous):

Just plug the points in

OpenStudy (anonymous):

I'm still confused...

OpenStudy (anonymous):

(-1,1) and (1,1) y = x2 1=(-1)^2 true y = –x 1=-(-1) true y = x2 1=(1)^2 true y = –x 1 = -1 false first set doesn't work

OpenStudy (anonymous):

Refer to the Mathematica attachment.

OpenStudy (anonymous):

(0,0) y = x2 0 = 0^2 true y = –x 0 = -0 true

OpenStudy (anonymous):

(0, 0) and (–1, 0) y = x2 0 = 0^2 true y = –x 0 = -0 true y = x2 0 = -1^2 false this set doesn't work y = –x

OpenStudy (anonymous):

So it would be (0,0). Right?

OpenStudy (anonymous):

From Oliveeeuuuh attachment...look where the lines intersect. That represents your solution

OpenStudy (anonymous):

heyyyyyyy

OpenStudy (anonymous):

(0,0) is only one of the solutions.